# Spherical Double K-Means: a co-clustering approach for text data analysis

### v1

@@ Abstract
In this study, we introduce the Spherical Double K-Means (SDKM) clustering method for text data. A novel approach for simultaneous clustering of terms and documents. Using the strengths of k-means, double k-means, and spherical k-means, SDKM addresses the challenges of high dimensionality, noise, and sparsity inherent in text analysis. We address the choice of the number of clusters, both for the words and documents, using the cluster validity index pseudo-$F$, and verify the reliability of the method through simulation studies.
@@ 1 Introduction
In recent years, the exponential growth of textual data from various sources such as social networks, electronic health records, insurance claims, and news outlets has significantly increased the importance of effective text analysis methods. Text clustering, which involves grouping similar documents without predefined categories or labels (Iezzi, 2012), plays a pivotal role in uncovering hidden patterns, topics, or clusters within large corpora. This facilitates the exploration and organization of vast text datasets based on inherent similarities, aiding in information retrieval, topic modeling, and data summarization.
Among the most used clustering algorithms, k-means and its variants are valued for their simplicity, scalability, and ability to quickly classify large and unstructured data. However, traditional k-means clustering has limitations when applied to text data. It requires the pre-specification of the number of clusters, is sensitive to outliers, and may converge to local minima. Moreover, text data often exhibit high dimensionality, sparsity, and noise, making the Euclidean distance metric used in standard k-means less suitable. This is because Euclidean distance can overemphasize longer documents with higher term weights, leading to biased clustering results (Hornik et al., 2012).
@@
The remainder of the paper is structured as follows: Section 2 provides background on text data processing and reviews relevant clustering methodologies. Section 3 introduces the SDKM methodology in detail. Section 4 presents the data generation procedure, the pseudo-$F$ index for the selection of the optimal number of clusters, and a simulation study to evaluate the performance of the proposed method. Section 5 showcases the application of SDKM to the US presidential inaugural addresses. Finally, Section 6 concludes the paper and discusses potential directions for future research.
@@ 2.1 Addressing text data
    Stemming or Lemmatization: Converting each token to its base form (e.g., “studies”, “studying”, “studied” to “study”), grouping morphological variants of the same word.
@@ 2.2.1 Double k-means: DKM
If the units’ membership matrix $\bm{U}$ degenerates into an identity matrix of order $K$, i.e. $\bm{U}=\mathbbm{1}_{K}$, the model 1 can be written as $\bm{X}=\bm{\overline{Y}V}^{\prime}+\bm{E}$, which is the model of the KM to partition variables into $Q$ clusters.
Similarly, if the variables’ membership matrix $\bm{V}$ degenerates into an identity matrix of order $J$, i.e. $\bm{V}=\mathbbm{1}_{J}$, the model 1 can be written as $\bm{X}=\bm{U\overline{Y}}+\bm{E}$, which is the model of the KM to partition units into $K$ clusters.
@@ 3 Methodology: Spherical double k-means
We integrate SKM in DKM in order to simultaneous cluster units and variable with spherical shape clusters. We therefore introduce the Spherical double k-means (SDKM). SDKM methodology aims to incorporate the advantages of SKM, i.e. its ability to address over-dispersion in the data and detect noise, into a double clustering scenario.
@@ 4.5 Performance in terms of true partitions and centroids recovery
ARI for U mean 1.000 1.000 1.000 0.974 0.907 0.794 0.675 0.587 0.480 0.354
median 1.000 1.000 1.000 1.000 0.976 0.877 0.682 0.567 0.467 0.360
ARI for V mean 1.000 1.000 1.000 0.984 0.933 0.840 0.717 0.649 0.507 0.332
median 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.866 0.522 0.192
@@ 5 Application
We applied the SDKM method on a dataset containing the US presidential inaugural addresses. These speeches were delivered by all US presidents from George Washington in 1789 to Joe Biden in 2021. We aim to investigate if any shared topics can be identified in clusters of speeches delivered by different presidents. The initial corpus comprises 59 documents, with a total of 151,536 tokens, 9,442 types, and 4,218 hapaxes. On average, each document contains 770.4 types and 2,568 tokens. We calculated the type/token ratio, which is 0.0623, and the hapax percentage, which is 0.4467. These values suggest a high lexical variety, justifying the computational analysis of the corpus. Specifically, we used the DSKM to cluster both the documents and the words simultaneously. During the first stage of our analysis, we focused on cleaning and pre-processing the data. This step is crucial for text analysis. To achieve this, we used tokenization to break the text into smaller units and removed numbers, punctuation, symbols, and separators. We also applied lemmatization to bring each token to its basic form. To create a document-term matrix, we removed stop-words and trimmed the data by removing all words with a frequency less than or equal to 11. This part of the analysis is fundamental in order to highlight the main topics treated by the presidents. We utilized the Tf-idf weighting scheme, which evaluates the significance of a word for a text or class of texts (Liang & Niu, 2022). Manning (2009). Tf-idf measures how important a term is within a document relative to a corpus, and it is calculated as:
$x_{ij}=\frac{n_{ij}}{n_{\cdot j}}\log_{10}\frac{M}{m}.$
Where $n_{ij}$ is the raw number of occurrences, $M$ is the number of documents and $m$ is the number of documents that include the term. Using the Tf-idf normalized term frequency, we aim to avoid any bias towards longer documents, which may have a higher term count even if the term is not particularly important. After pre-processing, we ended up with a final data matrix consisting of 59 documents and 1,206 types, this matrix results to be sparse at 67.73%.
@@ 5.1 Spherical Double k-means
To determine the best number of clusters for both terms and documents, we used the Pseudo-$F$ index with K and Q ranging from 2 to 10. The results are displayed in Table 4.
Table 4: Pseudo-F values for combinations of $K$ and $Q$
@@
2 1199.49 675.70 655.62 410.47 452.41 236.09 236.18 286.87 212.85
3 707.30 593.86 351.88 241.79 268.77 190.25 162.91 183.58 181.71
4 536.64 417.85 344.37 278.05 167.46 168.04 140.40 131.18 103.20
5 439.65 293.96 257.97 166.10 187.59 130.32 91.69 88.65 131.21
6 358.84 257.03 193.37 150.41 153.65 142.23 99.68 62.87 75.54
7 326.77 190.50 172.07 134.73 90.79 91.78 91.05 59.24 57.11
8 273.30 192.00 162.46 126.16 139.08 84.05 76.20 76.54 67.62
9 246.95 167.39 116.41 107.27 91.55 73.29 64.11 64.56 50.78
10 208.50 165.17 127.04 98.14 82.05 62.77 70.39 46.00 50.75
The highest Pseudo-$F$ value occurs at $K=2$ and $Q=2$, suggesting that the optimal number of clusters for both words and documents is 2. However, as discussed in Section 4.3, the Pseudo-$F$ index can tend to underestimate the number of clusters. The second highest Pseudo-$F$ value occurs at $K=3$ and $Q=2$, and we found that this combination provides much more interpretable results. Interpretability is a key consideration in cluster analysis (Fraley & Raftery, 1998), and therefore we decided to use $K=3$ and $Q=2$ for our analysis.
The analysis provides insights into the evolution of themes in U.S. presidential inaugural addresses. As shown in Figure 2, the three clusters of words contain 457, 420 and 328 words, while Figure 3 shows that the two groups of documents contain 31 and 28 documents.
Figure 2: Number of words by cluster
Figure 3: Number of documents by cluster
Figure 4 illustrates how the documents are distributed across the clusters identified with the SDKM, along with some of the main events in U.S. history. It is evident that the two groups have a chronological distinction between them. Cluster 1 encompasses inaugural speeches from 1789 to 1861, reflecting the early years of the nation, while Cluster 2 includes speeches from 1865 onwards, representing modern presidential rhetoric.
Figure 4: Chronologically ordered US presidential inaugural addresses, colored by cluster of documents
@@
Figure 5 shows the 30 words with the highest frequencies by cluster of words.
Figure 5: Most frequent words by cluster of words
@@
Analyzing the most frequently used words within each cluster allowed us to identify distinct themes. Cluster 1, which we have termed the “American Dream” cluster, includes words such as “america”, “freedom”, “democracy”, and “dream”. These words suggest a focus on patriotism, aspiration, and forward-looking ideals. Cluster 2, labeled “Law and Order”, contains terms like “union”, “constitution”, “state”, and “government” indicating an emphasis on legal structures and national unity. Finally, cluster 3, referred to as “Politics, Economics, and Secession” features words such as “congress”, “law”, “business”, and “policy”, pointing to themes of political processes and economic considerations.
Figure 6 displays the 30 most frequent words within the two clusters of documents, colored by their respective word clusters. The most frequent words in the second cluster of documents all belong to word cluster 1. In contrast, the first cluster of documents includes frequent words mainly from word cluster 2, with some from cluster 3, particularly those related to politics and economic matters.
@@
Figure 6: Most frequent words in each cluster of documents, colored by cluster of words
Cluster 1 documents predominantly use words from word clusters 2 and 3, reflecting the nation’s focus during its formative years on establishing governance, constitutional foundations, and addressing issues like secession. On the other hand, cluster 2 documents largely utilize words from word cluster 1, highlighting a shift towards aspirational themes, national identity, and the “American dream” in modern times. This evolution of presidential rhetoric mirrors the nation’s historical trajectory, transitioning from foundational concerns to aspirations for the future. Significant historical events, such as the civil war, world wars, and the great depression, have influenced the themes and language of inaugural addresses. The clustering results underscore how presidents adapt their messages to address the pressing issues of their times while shaping the nation’s collective vision.
6 Final consideration

### v2

@@ Abstract
In text analysis, Spherical K-means (SKM) is a specialized k-means clustering algorithm widely utilized for grouping documents represented in high-dimensional, sparse term-document matrices, often normalized using techniques like TF-IDF. Researchers frequently seek to cluster not only documents but also the terms associated with them into coherent groups. To address this dual clustering requirement, we introduce Spherical Double K-Means (SDKM), a novel methodology that simultaneously clusters documents and terms. This approach offers several advantages: first, by integrating the clustering of documents and terms, SDKM provides deeper insights into the relationships between content and vocabulary, enabling more effective topic identification and keyword extraction. Additionally, the two-level clustering assists in understanding both overarching themes and specific terminologies within document clusters, enhancing interpretability. SDKM effectively handles the high dimensionality and sparsity inherent in text data by utilizing cosine similarity, leading to improved computational efficiency. Moreover, the method captures dynamic changes in thematic content over time, making it well-suited for applications in rapidly evolving fields. Ultimately, SDKM presents a comprehensive framework for advancing text mining efforts, facilitating the uncovering of nuanced patterns and structures that are critical for robust data analysis.
@@ 1 Introduction
In recent years, the rapid increase in textual data from various sources, such as social networks, electronic health records, insurance claims, and news outlets, has made effective text analysis methods more critical than ever. Text clustering, the process of grouping similar texts, is crucial in this analysis. Text clustering, which involves grouping similar documents without predefined categories or labels (Iezzi, 2012), plays a pivotal role in uncovering hidden patterns, topics, or clusters within large corpora. This facilitates the exploration and organization of vast text datasets based on inherent similarities, aiding in information retrieval, topic modeling, and data summarization.
Among the most used clustering algorithms, k-means and its variants are valued for their simplicity, scalability, and ability to classify large and unstructured data quickly. However, traditional k-means clustering has limitations when applied to text data. The algorithm’s reliance on Euclidean distance often fails to accurately represent relationships in high-dimensional spaces, such as those created by text representations like TF-IDF or bag-of-words, ignoring meaningful semantic relationships like synonyms and context. Additionally, the high dimensionality of text data can render Euclidean distance less meaningful, leading to ineffective clustering results. Another challenge is k-means’ sensitivity to initialisation, as its performance heavily depends on the number of clusters and the initial placement of centroids, which can result in unstable or suboptimal outcomes. The algorithm also assumes that clusters are spherical and evenly distributed, a condition rarely met in text data, which often forms non-convex or overlapping clusters. Furthermore, k-means treats text as numerical data, lacking semantic understanding, which can lead to similar words being assigned to different clusters. Lastly, k-means is sensitive to outliers and noise, such as rare words or misspellings, distorting cluster centroids and reducing overall clustering quality. Moreover, text data often exhibit high dimensionality, sparsity, and noise, making the Euclidean distance metric used in standard k-means less suitable. This is because Euclidean distance can overemphasize longer documents with higher term weights, leading to biased clustering results (Hornik et al., 2012).
@@
The remainder of the paper is structured as follows: Section 2 provides background on text data processing and reviews relevant clustering methodologies. Section 3 introduces the SDKM methodology in detail. Section 4 presents the data generation procedure, the pseudo-$F$ index for the selection of the optimal number of clusters, and a simulation study to evaluate the performance of the proposed method. Section 5 showcases the application of SDKM to the US presidential inaugural addresses. In Section 6, we compare the results achieved with SDKM to those achieved with Double K-Means (DKM) in the proposed application, highlighting their similarities and differences. Finally, Section 7 concludes the paper and discusses potential directions for future research.
@@ 2.1 Addressing text data
    Stemming or Lemmatization: Stemming reduces words to their root or base form by removing suffixes or prefixes, often using heuristic rules. Lemmatization reduces words to their canonical form (lemma) based on dictionary definitions and linguistic analysis. For example, “studies”, “studying”, “studied” are reduced to “study”.
@@ 2.2.1 Double k-means: DKM
If the units’ membership matrix $\bm{U}$ degenerates into an identity matrix of order $K$, i.e. $\bm{U}=\bm{1}_{K}$, the model 1 can be written as $\bm{X}=\bm{\overline{Y}V}^{\prime}+\bm{E}$, which is the model of the KM to partition variables into $Q$ clusters.
Similarly, if the variables’ membership matrix $\bm{V}$ degenerates into an identity matrix of order $J$, i.e. $\bm{V}=\bm{1}_{J}$, the model 1 can be written as $\bm{X}=\bm{U\overline{Y}}+\bm{E}$, which is the model of the KM to partition units into $K$ clusters.
@@ 3 Methodology: Spherical double k-means
We integrate SKM in DKM to simultaneously cluster units and variables with spherical shape clusters. We therefore introduce the Spherical double k-means (SDKM). SDKM methodology aims to incorporate the advantages of SKM, i.e. its ability to address over-dispersion in the data and detect noise, into a double clustering scenario.
@@ 4.5 Performance in terms of true partitions and centroids recovery
ARI for U mean 1 1 1 0.974 0.907 0.794 0.675 0.587 0.480 0.354
median 1 1 1 1 0.976 0.877 0.682 0.567 0.467 0.360
ARI for V mean 1 1 1 0.984 0.933 0.840 0.717 0.649 0.507 0.332
median 1 1 1 1 1 1 1 0.866 0.522 0.192
@@ 5 Application
We applied the SDKM method to a dataset containing U.S. presidential inaugural addresses, ranging from those delivered by George Washington in 1789 to Joe Biden in 2021. Our objective is to identify shared topics among speeches given by different presidents. Using SDKM allows for the simultaneous clustering of both documents and terms, providing a comprehensive view of the evolving presidential discourse. The initial corpus comprises 59 documents, with 151,536 tokens, 9,442 types, and 4,218 hapaxes. On average, each document contains 770.4 types and 2,568 tokens. We calculated the type/token ratio, 0.0623, and the hapax percentage, 0.4467. These values suggest a high lexical variety, justifying the computational analysis of the corpus.
In our analysis, data cleaning and preprocessing were crucial. This involved tokenizing texts and removing extraneous elements such as numbers and punctuation. Lemmatization was applied to reduce each word to its base form. We built a document-term matrix, excluding stop-words and trimming infrequent words with a frequency of 11 or less. These steps were necessary to focus the analysis on core topics articulated by the presidents. We employed the TF-IDF weighting scheme, which assesses a word’s significance within documents relative to the entire corpus ((Liang & Niu, 2022), Manning (2009)). This approach helps mitigate length-induced bias in document analysis by focusing on term importance rather than sheer frequency. Specifically, TF-IDF measures how important a term is within a document relative to a corpus, and it is calculated as:
$x_{ij}=\frac{n_{ij}}{n_{\cdot j}}\log_{10}\frac{M}{m},$
where $n_{ij}$ is the raw number of occurrences, $M$ is the number of documents and $m$ is the number of documents that include the term. The final data matrix, comprising 59 documents and 1,206 types, demonstrated a sparsity of 67.73%
@@ 5.1 Spherical Double k-means
To determine the best number of clusters for both terms and documents, we used the Pseudo-$F$ index with K and Q ranging from 2 to 10. Table 4 presents these Pseudo-F values, highlighting the clustering dynamics.
Table 4: Pseudo-F values for combinations of $K$ and $Q$.
@@
2 1199.5 675.7 655.6 410.5 452.4 236.1 236.2 286.9 212.9
3 707.3 593.9 351.9 241.8 268.8 190.3 162.9 183.6 181.7
4 536.6 417.9 344.4 278.1 167.5 168.0 140.4 131.2 103.2
5 439.7 294.0 258.0 166.1 187.6 130.3 91.7 88.7 131.2
6 358.8 257.0 193.4 150.4 153.7 142.2 99.7 62.9 75.5
7 326.8 190.5 172.1 134.7 90.8 91.8 91.0 59.2 57.1
8 273.3 192.0 162.5 126.2 139.1 84.1 76.2 76.5 67.6
9 247.0 167.4 116.4 107.3 91.6 73.3 64.1 64.6 50.8
10 208.5 165.2 127.0 98.1 82.1 62.8 70.4 46.0 50.8
The highest Pseudo-$F$ value occurs at $K=2$ and $Q=2$, suggesting that the optimal number of clusters for both words and documents is 2. However, as discussed in Section 4.3, the Pseudo-$F$ index can tend to underestimate the number of clusters. The second highest Pseudo-$F$ value occurs at $K=3$ and $Q=2$. The observed patterns in the clusters indicate that the choice of $K$ = 3 and $Q$ = 2 produces clusters with greater interpretability, a crucial factor in the effective analysis of clusters as fully discussed in (Fraley & Raftery, 1998). In addition, to further investigate this aspect, we apply our methodology with $K=10$ and $Q=10$ and we plot the clustergram of the centroid matrix. The clustergram in Figure 2, which shows the hierarchical clustering of the rows and columns of the centroids matrix through a heatmap and two dendrograms, reveals $K=3$ distinct clusters of words and $Q=2$ distinct clusters of documents. The dendrograms are obtained by using one minus the sample correlation between points as distance metric for both rows and columns and single linkage and Ward linkage as hierarchical clustering methods, for rows and columns, respectively. In addition, the same clustergram procedure is applied to a sample of $150$ rows of the theoretical data matrix reconstructed via: $\mathbf{X}_{t}=\mathbf{U}\bar{\mathbf{Y}}\mathbf{V}^{\prime}$. Figure 3 shows again $K=3$ distinct clusters of words and $Q=2$ distinct clusters of documents.
Figure 2: Clustergram of $10\times 10$ centroids matrix.
Figure 3: Clustergram of the theoretical data matrix narrowed down to 150 sampled rows.
Therefore, $K$ = 3 and $Q$ = 2 was our final choice.
The analysis provides insights into the evolution of themes in U.S. presidential inaugural addresses. The three clusters of words contain 457 (37.9%), 420 (34.9%) and 328 (27.2%) words, while the two groups of documents contain 31 and 28 documents.
5.1.1 Document Distribution and Historical Context
Figure 4 presents the document distribution across SDKM’s two clusters, along with key events in U.S. history. Cluster 1 spans speeches from 1789 to 1861, reflecting the early nation-building era, while Cluster 2 covers addresses from 1865 onward, aligning with the modern evolution of presidential rhetoric.
Figure 4: Historical timeline of the US presidents, colored by cluster of documents and their 5 most frequent words (TF-IDF normalized).
@@ 5.1.1 Document Distribution and Historical Context
5.1.2 Word Clusters and Thematic Interpretation
@@
Figure 5 displays the top 30 terms in each of the three word clusters. Cluster 1, that we labeled “American Dream,” features words like “america,” “freedom,” “democracy,” and “dream,” pointing to aspirational rhetoric and national ideals. Cluster 2 has been labeled “Law and Order,” it centers on terms like “union,” “constitution,” “state,” and “government,” underscoring legal and institutional frameworks. We called Cluster 3, “Politics, Economics, and Secession.” It includes words such as “congress,” “law,” “business,” and “policy,” reflecting deeper political processes and economic discussions.
Figure 5: Top 30 terms in each of the three word clusters (TF-IDF weighted).
Figure 6 further illustrates the document-level clustering in a pseudo-3D coordinate system, where each axis corresponds to a cluster of words. Notably, speeches in the “Politics, Economics, and Secession” dimension align more strongly with the older cluster, while documents with higher “American Dream” or “Law and Order” components fall into the more modern cluster. Addresses like Lincoln’s 1865 or Coolidge’s 1925 appear in positions consistent with their heavier emphasis on unity or policy-related concerns, linking them more closely to modern rhetorical changes.
Figure 6: Each point represents a single inaugural address, colored by its document-cluster membership (blue for cluster 1, green for cluster 2). The coordinates are mean TF-IDF values of each address in the three word-clusters.
Finally, Figure 7 examines the 30 most frequent words in each document cluster, colored by word-cluster membership. The second document cluster draws heavily on the “American Dream” word set, whereas the first cluster incorporates terms largely from “Law and Order,” as well as some from “Politics, Economics, and Secession.” This pattern aligns with the nation’s historical development: early speeches concerned with legal underpinnings and secession disputes, and later addresses emphasizing aspirational, forward-facing themes.
@@
Figure 7: Most frequent words in each document cluster, colored by their respective word clusters.
Altogether, these results highlight how the themes covered by the presidents’ addresses evolved in time and how major historical events, such as the Civil War, world wars, and the Great Depression, shaped the language and focus of presidential inaugural addresses, illustrating a broader shift from foundational legal concerns to modern themes of national identity and aspirations for the future.
6 Comparison of DKM and SDKM Clusters
We applied Double K-Means (DKM) on the same data (preprocessed and TF-IDF weighted) to compare the results with those achieved using SDKM. We examine both document-level and word-level clusters, highlighting how Euclidean distance (DKM) versus cosine similarity (SDKM) can produce subtle but meaningful differences in this double clustering setting. We do not dwell on the comparison of computational times, as they are negligible in both methods.
6.1 Document Clusters: Lincoln 1865 as the Only Difference
Regarding the documents cluster membership, strikingly, every speech falls into the same two‐cluster partition across both methods, earlier historical presidents vs. later ones, except for Lincoln’s 1865 address.
  • 1.
    DKM: Places Lincoln’s 1865 inaugural in the older (founding era) cluster. Even with TF-IDF weighting, Euclidean distance reflects the overall magnitude of the feature vector; Lincoln’s 1865 speech aligns more closely with earlier addresses in that sense.
  • 2.
    SDKM: Classifies Lincoln 1865 among more modern addresses. By focusing on relative usage proportions of words like “union,” “heal,” “mercy,” and “nation,” SDKM finds it more thematically aligned with 20th‐century rhetoric (rather than strictly early 19th‐century addresses).
6.2 Word Clusters: Thematic Comparisons
Beyond their broad similarity in document partitioning, DKM and SDKM exhibit noteworthy differences at the word level. To illustrate how each method groups the main words, we identify the most frequent terms in each cluster, calculate pairwise Jaccard distances among these word sets, and then apply classical multidimensional scaling (MDS).
Label the three DKM clusters as DKM_1, DKM_2, DKM_3 and the three SDKM clusters as SDKM_1, SDKM_2, SDKM_3. For each cluster, we measure the overlap of the main words using the Jaccard distance:
$d_{\text{Jaccard}}(A,B)\;=\;1\;-\;\frac{|A\cap B|}{|A\cup B|},$
where $A$ and $B$ are sets of top words. A smaller value indicates stronger overlap in membership. The resulting $3\times 3$ dissimilarity matrix appears in Table 5.
Table 5: Jaccard dissimilarities among the three clusters of DKM (rows) and SDKM (columns). Lower values imply greater overlap of top words.
SDKM_1 SDKM_2 SDKM_3
DKM_1 0.095 1 0.947
DKM_2 1 0.462 0.824
DKM_3 0.974 1 0.947
Figure 8 then shows the MDS projection of these distances into a 2D space. Clusters that share more words appear closer, while those sharing fewer words appear farther apart. One cluster, centered on words such as “america,” “freedom,” “democracy,” and “dream,”, emerges as quite similar in both methods, whereas the other two clusters diverge more substantially.
Figure 8: MDS plot of word clusters for DKM (green) and SDKM (orange).
Both DKM and SDKM isolate a cluster of aspirational or national-identity terms, including “america,” “freedom,” “democracy,” “child,” “peace,” and “dream.” These words typically co-occur in modern addresses (post Civil War and throughout the 20th century) emphasizing a forward-looking vision for the nation. Because such vocabulary dominates many speeches from that era, Euclidean and cosine-based approaches alike identify a similar “American Dream” cluster.
The main discrepancies arise in how each method groups governance, policy, and constitutional terms. In the SDKM application, for instance, older addresses (1789–1861) focus heavily on “union,” “constitution,” “secession,” and “law,” while later ones shift toward “dream,” “freedom,” “economics,” and “America’s ideal”. Although DKM also detects an older vs. newer thematic line, it often merges older constitutional references with modern economic terms if they appear together in certain large documents. By contrast, SDKM, emphasizing relative usage, splits them if they differ in proportional frequency. For example, words like “race,” “tax,” and “business” are grouped together by DKM if they frequently co-occur in a handful of lengthy speeches, whereas SDKM may place them in separate clusters if they appear in different ratios across those speeches.
These patterns underscore how SDKM’s cosine-based metric isolates specific rhetorical proportions, creating distinct clusters for governance/constitutional issues versus modern policy/economic themes. Meanwhile, DKM lumps some topics together whenever they coexist in the same documents, downplaying subtle proportional distinctions. This aligns with the document-level discrepancy observed in Lincoln’s 1865 speech: a borderline case, rich in “union” and “heal” language, can shift clusters under SDKM, just as certain policy or war-related terms shift between the two methods depending on how they recur in the corpus.
7 Final considerations